OpenStudy (radvedder):

11th grade math - graphing linear equations. help!!! determine whether each equation is a linear equation. if so, write the equation in standard form. y=2-3x 5x=y-4

OpenStudy (anonymous):

standard form looks like \[ax+by=c\]

OpenStudy (anonymous):

for the first one, add \(3x\) to both sides and you get it in one step

OpenStudy (radvedder):

yes, but i'm not sure how to get it like that. can you help me on a few of them?

OpenStudy (anonymous):

\[y=2-3x\\ 3x+y=2\] finito

OpenStudy (radvedder):

i'm not sure if i subtract 2 or add 3x?

OpenStudy (radvedder):

how'd you get it like that

OpenStudy (anonymous):

you want all the variables on the left, so you would add \(3x\) to both sides

OpenStudy (anonymous):

by "add" i just mean write it

OpenStudy (radvedder):

ohh i see!

OpenStudy (diamondboy):

y= mx + b is d standard form for linear equations

OpenStudy (radvedder):

is it still not in standard form? 3x+y=2

OpenStudy (anonymous):



OpenStudy (radvedder):

yes, i have 3x+y=2 now. do i leave it at that?

OpenStudy (anonymous):

yes \[3x+y=2\] is STANDARD FORM

OpenStudy (anonymous):

leave it be, don't mess with it

OpenStudy (diamondboy):

for linear equations?

OpenStudy (anonymous):

@diamondboy yes, \[ax+by=c\] standard form \[y=mx+b\] slope intercept form

OpenStudy (diamondboy):

yep u are right I am so sorry

OpenStudy (radvedder):

and the next one is 5x=y-4

OpenStudy (radvedder):

i would add 4?

OpenStudy (anonymous):

no

OpenStudy (diamondboy):

subtract y from both sides

OpenStudy (anonymous):

you want the variables on the left subtract \(y\)

OpenStudy (radvedder):

5x-y=4

OpenStudy (anonymous):

no

OpenStudy (radvedder):

which is standard, yes?

OpenStudy (anonymous):

the \(-4\) is what stays on the right

OpenStudy (anonymous):

\[5x=y-4\\ 5x-y=-4\] the minus sign does not go away

OpenStudy (radvedder):

oh, i had that. i just typed it to you wrong

OpenStudy (radvedder):

and would xy=6 already be in standard form